A magnetic circuit consists of a cast steel yoke which has a cross-sectional area of 200 mm2 and a mean length of 120 mm. There are two air gaps, each 0.2 mm long. Calculate the mmf required to produce a

flux of 0.5 mWb in the air gaps and the value of the relative permeability of cast steel at this flux density. The magnetization curve for cast steel is given by the following:

B (T) 0.1 0.2 0.3 0.4

H (A/m) 170 300 380 460


1
Expert's answer
2020-06-29T14:19:32-0400

As per the given question,

Cross sectional area "(A)= 200mm^2"

Mean length "(l)=120mm"

Length of air gap=0.2mm

Magnetic flux "(\phi)=0.5 mWb"

"B =0.1"

"H =170"

We know that "\mu_r\mu_o=\frac{B}{H}"

"\Rightarrow F=\frac{\phi l}{\mu_r \mu A}"


"\Rightarrow F=\frac{0.5 mWb\times 120mm}{\frac{B}{H}\times200mm^2}"

"\Rightarrow F=\frac{60\times 170}{0.1}=102000N"